OpenStudy (anonymous):

Medal will be awarded! Algebra 1

jimthompson5910 (jim_thompson5910):

Use the formula SA = 2*pi*r^2 + 2*pi*r*h where SA = surface area r = radius h = height and you can use 3.14 for pi

OpenStudy (anonymous):

I got 803.84 So the answer is 803.8

jimthompson5910 (jim_thompson5910):

one sec while I check

OpenStudy (anonymous):

Here is another problem that relates to THE CYLINDER Image: 4. Find the volume of the cylinder. (1 point) 1607.7 in.^3 2,845.7 in.^3 6,430.7 in.^3 401.9 in.^3

jimthompson5910 (jim_thompson5910):

803.8 is correct let me look at the other one real quick

jimthompson5910 (jim_thompson5910):

now you use the formula V = pi*r^2*h and you plug in r = 8, h = 8 and replace pi with 3.14 V = 3.14*8^2*8 V = 1,607.68 V = 1,607.7

jimthompson5910 (jim_thompson5910):

so looks like choice A

OpenStudy (anonymous):

Okay thank you.

jimthompson5910 (jim_thompson5910):

yw